# Leveraging Large Language Models for Explainable Activity Recognition in Smart Homes: A Critical Evaluation

### v1

@@ Abstract.
This manuscript is under review. Please refer to authors for up-to-date information.
@@ 1. Introduction
In this paper, we critically evaluate the role that LLMs may potentially have in explainable sensor-based HAR (XAR). There have been preliminary investigations of applying LLMs to HAR (Cleland et al., 2024), leveraging the common-sense knowledge about human activities encoded in LLMs to reduce the required labeled data to train supervised classifiers. However, to the best of our knowledge, no existing work studies how LLMs can also be leveraged for explanations and the possible dangers they may introduce. We aim to answer these research questions:
@@
Q3) What are the drawbacks and risks in using LLM for XAR?
Our results answer positively to Q1 by implementing an end-to-end LLM-based system and testing it on two public datasets, comparing recognition accuracy with a deep learning baseline and an LLM method that does not provide explanations. We also positively answer to Q2 by proposing a standard format that different XAI methods may adopt to represent the most relevant features leading to the prediction. By using this input and an appropriate prompting technique, LLMs can generate appreciated explanations as proved by user surveys. Regarding Q3, we discuss the potential disadvantages of such systems and consider possible mitigation actions. We identify over-reliance as the main risk, by providing examples of intuitive LLM explanations for wrong predictions. We also discuss cost and privacy issues, real-life deployment concerns, and the possible impact of LLMs’ hallucinations.
@@
  • While the results are promising, our critical evaluation highlights the risks and the drawbacks of adopting LLMs in XAR, including examples illustrating the risk of over-reliance on explanations.
@@ 2.2. Large Language Models for Human Activity Recognition
The HAR community is actively working on leveraging Large Language Models (LLMs). Several existing approaches take advantage of the intrinsic knowledge of LLMs about human activities to reduce the amount of required training data (Liu et al., 2023; Okita et al., 2023; Zhou et al., 2023; Zilelioglu et al., 2023; Arrotta et al., 2024; Thukral et al., 2024; Leng et al., 2024). Most interestingly, some recent works (most of them are still pre-prints) considered the more challenging setting of using LLMs for zero-shot recognition of human activities (Ji et al., 2024; Hota et al., 2024). These works leverage a textual representation of input sensor data to query the LLM for the activity most likely performed by the monitored subject. Several research papers recently proposed LLM-based zero-shot approaches for ADL recognition in smart homes, proposing specifically designed prompts enabling LLMs to reason on textual representation of input sensor data to identify the most likely ADLs performed by the inhabitants (Gao et al., 2024; Cleland et al., 2024; Xia et al., 2023; Civitarese et al., 2024; Chen et al., 2024; Fritsch et al., 2024). However, all of these works focus only on recognizing the performed activities, without providing explanations that non-expert users may leverage to grasp the rationale behind the decision.
@@ 3.1. The Explainable Human Activity Recognition Problem
We first formally define the XAR problem that we aim to address in this work. Like many studies in this field, we focus on a smart home setting where either a single individual resides or the sensing system is capable of accurately linking each sensor activation to the specific subject responsible for triggering it. The home is equipped with several environmental sensors that capture the interaction of the subject with the surrounding environment and the person may be carrying a wearable device revealing low-level physical activities. Depending on its type, each sensor generates a data stream corresponding to activation/deactivation (e.g., open/close), measurements (e.g., humidity values), or inertial sensor data. Measurements and inertial data are preprocessed to generate higher-level information (e.g., high humidity or sitting posture) so that each sensor has an associated time series of events that identify the start and end time of the detected conditions (e.g., high humidity start/end, sitting start/end). The multi-variate time series corresponding to the whole set of sensors is segmented in temporal windows.
Given a window $\mathbf{w}$ in the time series of sensor events, and a set of candidate activities $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$, the goal of HAR is to predict the most likely activity $\hat{A}\in\mathcal{A}$ that the subject is performing during the time of that window. This task is usually performed by a classifier $h$ such that $h\left(\mathbf{w}\right)=\hat{A}$. Explainable classifiers generally associate a value to each event in $\mathbf{w}$, denoting its importance for the prediction $h\left(\mathbf{w}\right)$. For each prediction $\hat{A}=h\left(\mathbf{w}\right)$, an explanation $\mathbf{e}(h\left(\mathbf{w}\right))$ is defined as the set of events in $\mathbf{w}$ considered by $h$ as the most important for its prediction (e.g., based on a threshold on the importance values).
@@ 3.2. LLMe2e: Zero-Shot Explainable ADL Recognition
Starting from the work initially proposed in (Civitarese et al., 2024), we designed a novel Zero-Shot Explainable ADL recognition method based on LLMs. We will refer to this approach as LLMe2e. This method adopts an end-to-end approach, starting with sensor data and using a single LLM prompt to perform both ADL classification and natural language explanation generation simultaneously. LLMe2e does not require any training data, since it only leverages the intrinsic knowledge about human activities encoded in LLMs. The main differences with the work in (Civitarese et al., 2024) are a different representation of the input sensor data and the generation of explanations for each prediction. Note that even other zero-shot approaches proposed in the literature (e.g.,  (Cleland et al., 2024)) could be similarly extended to generate explanations. The high-level architecture of LLMe2e is shown in Figure 1.
@@
First, the raw stream of sensor data is transformed into a higher-level representation, referred to as sensor states, by the Sensor States Generation module. Intuitively, each sensor state represents the time interval during which a specific property detected by a sensor (e.g., the front door being opened) is true.
Next, the stream of sensor states is divided into fixed-time windows by the Segmentation module. Each window is then converted into a structured text format by the States2JSON module. Figure 2 shows a JSON standard format that we designed for sensor state representation. This format is used to provide the input to LLMe2e for activity prediction, but is also adopted by the LLMExplainer approach to represent the states corresponding to important events identified by XAR techniques, as we will discuss in Section 3.3.
@@
3.2.1. Sensor State Generation
[…]
3.2.2. Segmentation
The Segmentation module segments the stream of sensor states into fixed-length time windows of $\tau$ seconds, with an overlap factor of $o$ between consecutive windows. Each time window is associated with the sensor states that are active during that period, but only considering the overlapping portions of the original sensor state intervals.
@@ 3.2.3. System Prompt
3.2.4. User Prompt Generation
Each user prompt defines the specific instance of the problem that the LLM has to process. In LLMe2e, the User Prompt Generation generates a user prompt by leveraging the JSON representation of the input window created by the States2JSON module, also appending instructions to adopt the well-known Chain of Thought (CoT) prompting strategy (Wei et al., 2022). An example of a user prompt is shown in Figure 4.
@@ 3.3. LLMExplainer: LLMs as explanation generators for data-driven XAR
In this Section, we propose LLMExplainer: an LLM-based method to generate natural language explanations from the most important events $\mathbf{e}$ derived by a data-driven XAR methods. LLMExplainer is agnostic to the underlying XAR methods being adopted, since the only requirement is that their output, i.e., the most important events contributing to each prediction, can be represented in the JSON format depicted in Figure 2. Considering the existing XAR methods, the conversion in this format is straightforward. The architecture of LLMExplainer is depicted in Figure 6.
@@ 3.3.2. User Prompt
As mentioned above, LLMExplainer can be easily used with any XAR model. However, an additional module converting the most important features derived by a specific XAR model into the JSON format is required. Indeed, LLMExplainer expects important features expressed as states. Each state should be associated with a name (e.g., “the stove is on”), and the list of intervals in which that state was active. In order to clarify how LLMExplainer generates the user prompt, we show an example using DeXAR (Arrotta et al., 2022a), representing a state-of-the-art XAR method. The application of LLMExplainer to other XAR approaches like the one presented in (Das et al., 2023) is straightforward. Figure 10 shows an explanation (i.e., a heatmap) obtained by DeXAR’s classifier when predicting the activity preparing hot meal.
@@ 4.1. Datasets
For the sake of this work, we selected two publicly available smart-home datasets that include several types of environmental sensors, with one of them including also inertial sensors from a smartwatch. We excluded datasets based almost entirely on PIR sensors, like CASAS (Cook et al., 2009) for reasons that are discussed in Section 5.
@@ 4.1.2. MARBLE
The MARBLE (Arrotta et al., 2021) dataset includes data from $12$ subjects in both single and multi-inhabitant scenarios. For the sake of this work, we only consider the single-inhabitant scenarios of the dataset. The data was collected using smart plugs (to monitor stove and TV usage), magnetic (to monitor the interaction with drawers and cabinets), and pressure sensors (to monitor the usage of dining room chairs and the couch in the living room). Additionally, each subject wears a smartwatch to capture inertial data generated by hand gestures. Consistently with the work in (Arrotta et al., 2022a), we only differentiate between static and dynamic gestures. To detect incoming and outgoing phone calls, an Android application was installed on the smartphone of each subject. The original dataset includes 13 activities, but, following (Civitarese et al., 2024), we consider the following $11$ activities: clearing table, eating, entering home, leaving home, phone call, preparing cold meal, preparing hot meal, setting up table, taking medicines, using pc, and watching tv.
@@ 4.2. Baselines
  • ADL-LLM (Civitarese et al., 2024): a recent zero-shot ADL recognition method based on LLMs. We consider this method as a baseline for the recognition rate, as our LLMe2e approach is an extension and adaptation of it. Differently from LLMe2e, each window is translated into a sentence in natural language with a heuristic approach, rather than providing just the information of sensors’ triggering events like in the JSON template used in LLMe2e. ADL-LLM does not explicitly generate explanations, but we consider it to compare the recognition rate.
@@ 4.3. Experimental Setup
We developed the prototypes of LLMe2e and LLMExplainer in Python, querying the OpenAI ‘gpt-4o’ LLM model through the official OpenAI library, setting the temperature parameter to 00.

### v2

@@ Abstract.
Paper accepted for publication in the ACM Transactions on Internet Of Things (TIOT) journal.
@@ 1. Introduction
In this paper, we provide a critical evaluation of the role that LLMs may potentially have in eXplainable sensor-based HAR (XAR). On the one hand, we investigate how LLMs could be used to associate explanations with activity predictions. On the other hand, we also analyze the risks associated with using LLMs for the generation of explanations. For example, LLMs produce text that appears plausible but can be inaccurate and biased, and it can potentially lead to over-reliance (Hicks et al., 2024).
There have been preliminary investigations of applying LLMs to HAR (Cleland et al., 2024), leveraging the common-sense knowledge about human activities encoded in LLMs to reduce the required labeled data to train supervised classifiers. However, to the best of our knowledge, no existing work studies how LLMs can also be leveraged for explanations and the possible dangers they may introduce. We aim to answer these research questions:
@@
Q3) What are the drawbacks and risks of using LLM for XAR?
Our results answer positively to Q1 by implementing an end-to-end LLM-based system and testing it on two public datasets, comparing recognition accuracy with a deep learning baseline and an LLM method that does not provide explanations. We also positively answer to Q2 by proposing a standard format that different XAI methods may adopt to represent the most relevant features leading to the prediction. By using this input and an appropriate prompting technique, LLMs can generate appreciated explanations as proved by user surveys. Regarding Q3, we dedicate the whole Section 5 to discuss the potential disadvantages of such systems. We identify over-reliance as the main risk, by providing examples of intuitive LLM explanations for wrong predictions, and propose several mitigation strategies. We also discuss cost and privacy issues, real-life deployment concerns, and the possible impact of LLMs’ hallucinations.
To the best of our knowledge, this is the first systematic and critical exploration of using current LLMs capabilities to generate explanations in the Human Activity Recognition (HAR) domain. Moreover, this work proposes a novel pipeline to transform raw sensor data into a structured textual representation for LLMs, also describing how to capture the many sensors that can be deployed in a smart-home environment.
@@
  • We provide a critical evaluation highlighting the many risks and drawbacks of adopting LLMs in XAR, including examples illustrating the risk of over-reliance on explanations, and proposing several mitigation strategies.
@@ 2.2. Large Language Models for Human Activity Recognition
The HAR community is actively working on leveraging Large Language Models (LLMs). Several existing approaches take advantage of the intrinsic knowledge of LLMs about human activities to reduce the amount of required training data, mostly using inertial sensors obtained from mobile and wearable devices (Liu et al., 2023; Okita et al., 2023; Zhou et al., 2023; Zilelioglu et al., 2023; Arrotta et al., 2024; Leng et al., 2024; Sharma et al., 2025; Haresamudram et al., 2025; Li et al., 2024; Chen et al., 2024a). Interestingly, some of them considered the challenging setting of using LLMs for zero-shot recognition of human activities (Ji et al., 2024; Hota et al., 2024). These works leverage a textual representation of input sensor data to query the LLM for the activity most likely performed by the monitored subject. Since these approaches directly operate on raw inertial sensor data to detect low-level activities, they can not be directly compared with our approach, leveraging different type of data collected in smart home environments to recognize high-level ADLs.
Notably, LLMs have also been proposed for smart home ADL recognition (Thukral et al., 2024; Bouchabou et al., 2021), with some recent research proposing zero-shot approaches for ADL recognition, specifically designing prompts enabling LLMs to “reason” on textual representation of input sensor data to identify the most likely ADLs performed by the inhabitants (Gao et al., 2024; Cleland et al., 2024; Xia et al., 2023; Civitarese et al., 2025; Chen et al., 2024b; Fritsch et al., 2024). However, all of these works focus only on recognizing the performed activities, without providing explanations that non-expert users may leverage to grasp the rationale behind the decision.
@@ 3.1. The Explainable Human Activity Recognition Problem
We first formally define the XAR problem that we aim to address in this work. Like many studies in this field, we focus on a smart home setting where either a single individual resides or the sensing system is capable of accurately linking each sensor activation to the specific subject responsible for triggering it.
Depending on its type, each sensor generates a stream of semantic events. A semantic event is a tuple $\langle e,st,ts\rangle$ where $e$ defines the involved entity (e.g., fridge door), $st$ the status (e.g., open/close), and $ts$ the timestamp at which the event occurred. For binary sensors, converting raw measurements into semantic events is straightforward. These sensors naturally produce discrete outputs indicating state changes corresponding to specific points in time. For instance, if the magnetic sensor $M_{1}$ on the fridge triggers the value $1$ at time $t$, we generate a semantic event $\langle\textit{FridgeDoor},\textit{Opened},t\rangle$. On the other hand, sensors generating continuous values (e.g., temperature, humidity, inertial sensors, mmWave sensors) require more advanced processing to extract meaningful events. Considering continuous environmental sensors (e.g., temperature, humidity), heuristic thresholds can be applied. For example, if at time $t_{1}$ the humidity rises above a given threshold, a semantic event $\langle\textit{HighHumidity},\textit{Start},t_{1}\rangle$ is generated. Similarly, when the value falls below the threshold at time $t_{2}$, the event $\langle\textit{HighHumidity},\textit{End},t_{2}\rangle$ is generated. Finally, for more complex continuous sensors (e.g., inertial sensors, mmWave sensors, audio sensors), raw data (possibly preprocessed) is fed to simple classifiers in charge of detecting simple context conditions. For instance, a posture detection classifier analyzing inertial sensors from a wearable device may recognize at time $t$ that the subject was sitting, thus generating the semantic event $\langle\textit{Sitting},\textit{Start},t\rangle$. The resulting multivariate time series corresponding to the whole set of sensors is segmented in temporal windows.
Given a window $\mathbf{w}$ in the time series of semantic events, and a set of candidate activities $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$, the goal of HAR is to predict the most likely activity $\hat{A}\in\mathcal{A}$ that the subject is performing during the time of that window. This task is usually performed by a classifier $h$ such that $h\left(\mathbf{w}\right)=\hat{A}$. Explainable classifiers generally associate a value to each event in $\mathbf{w}$, denoting its importance for the prediction $h\left(\mathbf{w}\right)$. For each prediction $\hat{A}=h\left(\mathbf{w}\right)$, an explanation $\mathbf{e}(h\left(\mathbf{w}\right))$ is defined as the set of events in $\mathbf{w}$ considered by $h$ as the most important for its prediction (e.g., based on a threshold on the importance values).
@@ 3.2. LLMe2e: Zero-Shot Explainable ADL Recognition
Starting from the work initially proposed in (Civitarese et al., 2025), we designed a novel Zero-Shot Explainable ADL recognition method based on LLMs. We will refer to this approach as LLMe2e. This method adopts an end-to-end approach, starting with sensor data and using a single LLM prompt to perform both ADL classification and natural language explanation generation simultaneously. LLMe2e does not require any training data, since it only leverages the intrinsic knowledge about human activities encoded in LLMs. The main differences with the work in (Civitarese et al., 2025) are a different representation of the input sensor data and the generation of explanations for each prediction. Note that even other zero-shot approaches proposed in the literature (e.g.,  (Cleland et al., 2024)) could be similarly extended to generate explanations. The high-level architecture of LLMe2e is shown in Figure 1.
@@
First, the stream of semantic events (see Section 3.1) is transformed into a stream of sensor states by the Sensor States Generation module. Intuitively, each sensor state represents the time interval during which a specific property detected by a sensor (e.g., the front door being opened) is true. Next, the stream of sensor states is divided into fixed-time windows by the Segmentation module. Each window is then converted into a structured text format by the States2JSON module. Figure 2 shows a JSON standard format that we designed for sensor state representation. This format is used to provide the input to LLMe2e for activity prediction, but is also adopted by the LLMExplainer approach to represent the states corresponding to important events identified by XAR techniques, as we will discuss in Section 3.3.
@@ 3.2.1. System Prompt
3.2.2. User Prompt
The user prompt defines a specific instance of the problem that the LLM has to process. In the following, we describe how LLMe2e creates each instance from the stream of semantic events introduced in Section 3.1.
Step 1: Sensor State Generation
First, LLMe2e converts the stream of semantic events into a stream of states. Each state represents a time interval during which a specific property holds continuously (e.g., the fridge door was open from $t_{1}$ to $t_{2}$). Formally, a semantic state is denoted as $\text{st}(\sigma,t_{1},t_{2})$, indicating that the property $\sigma$ holds from time $t_{1}$ to $t_{2}$. For instance, $st(\textit{MedicineDrawerOpen},\textit{10:18am},\textit{10:19am})$ represents the fact that the medicine drawer was open in the interval $\textit{10:18am},\textit{10:19am}]$. As another example, $st(\textit{OnTheCouch},\textit{7:45pm},\textit{9:12pm})$ represents the fact that the pressure sensor on the couch reveals that the subject was sitting there between 7:45pm to 9:12pm. To derive a state, we pair semantic events that are complementary (e.g., Open/Close, On/Off). Two semantic events $\langle e,s,t_{1}\rangle$ and $\langle e,s^{\prime},t_{2}\rangle$ are paired if $s$ and $s^{\prime}$ are complementary, $t_{1}<t_{2}$, and there are no other events from $e$ occurring between $t_{1}$ and $t_{2}$. For instance, the state $st(\textit{OnTheCouch},\textit{7:45pm},\textit{9:12pm})$ is obtained from the semantic events $\langle\textit{OnTheCouch},\textit{Start},\textit{7:45pm}\rangle$ and $\langle\textit{OnTheCouch},\textit{End},\textit{9:12pm}\rangle$, given that there are no other $OnTheCouch$ events in the interval $\textit{7:45pm},\textit{9:12pm}]$.
Step 2: Segmentation
The Segmentation module segments the stream of sensor states into fixed-length time windows of $\tau$ seconds, with an overlap factor of $o$ between consecutive windows. Each time window is associated with the sensor states that are active during that period, but only considering the overlapping portions of the original sensor state intervals. Indeed, we want to ensure that each state is entirely included within the time window.
For example, consider a time window $w$ in the interval $\textit{3:20pm},\textit{3:40pm}]$. The semantic state $st(\textit{FridgeOpened},\textit{3:34pm},\textit{3:35pm})$ would be entirely contained in $w$. On the other hand, consider the state $st(\textit{TelevisionOn},\textit{3:12pm},\textit{3:25pm})$. In this case, while the state begins before the beginning of $w$, we would include only the overlapping portion $st(TelevisionOn,3:20pm,3:35pm)$. This step is crucial because we want to provide the LLM only with the properties that hold during the time window. More formally, a semantic state $st(\sigma,t_{1},t_{2})$ is considered for inclusion in a window $w$ if $t_{1},t_{2}]\cap[t_{1}^{w},t_{2}^{w}]\neq\emptyset$ with $t_{1}^{w},t_{2}^{w}$] being the interval of $w$ (i.e., the state has some time instants in common with $w$). The actual state that is included is $st(\sigma,t_{1}^{\prime},t_{2}^{\prime})$ where $t_{1}^{\prime}=\max(t_{1},t_{1}^{w})$ and $t_{2}^{\prime}=\min(t_{2},t_{2}^{w})$.
Step 3: States2Json
For each window $w$ of semantic states, the States2Json module creates a JSON representation, structured as follows. The “Time window” field is represented as an array of two timestamps, formatted as ”HH:MM:SS”, marking the start and end of $w$. This field is crucial to provide temporal context to the LLM. Then, for each unique property $\sigma$ of the semantic states in $w$, the JSON contains a dedicated field named with a “human readable” label for the property. For each field, the value is a list of one or more time intervals, one for each occurrence of a state with that property in the window.
Step 4: User Prompt Generation
Finally, the user prompt is generated by combining the JSON representation of the input window created by the States2JSON module with instructions to adopt the well-known Chain of Thought (CoT) prompting strategy (Wei et al., 2022). An example of generated user prompt is shown in Figure 4.
@@ 3.3. LLMExplainer: LLMs as explanation generators for data-driven XAR
In this Section, we propose LLMExplainer: an LLM-based method to generate natural language explanations from the most important events $\mathbf{e}$ derived by a data-driven XAR methods. LLMExplainer is agnostic to the underlying XAR methods being adopted, since the only requirement is that their output, i.e., the most important features contributing to each prediction, is represented as the most important semantic states contributing to classification. Indeed, LLMExplainer leverages the States2Json model described in Section 3.2.2 to convert semantic states into a JSON representation for the LLM. Considering the existing XAR methods, the conversion in this format is straightforward. The architecture of LLMExplainer is depicted in Figure 6.
@@ 3.3.2. User Prompt
As mentioned above, LLMExplainer can be easily used with any XAR model. However, an additional module converting the most important features derived by a specific XAR model into semantic states is required. Indeed, States2JSON expects important features expressed as states. Each state should be associated with a name (e.g., “the stove is on”), and the list of intervals in which that state was active. In order to clarify how LLMExplainer generates the user prompt, we show an example using DeXAR (Arrotta et al., 2022a), representing a state-of-the-art XAR method. The application of LLMExplainer to other XAR approaches like the one presented in (Das et al., 2023) is straightforward. Figure 10 shows an explanation (i.e., a heatmap) obtained by DeXAR’s classifier when predicting the activity preparing hot meal.
@@ 4.1. Datasets
While validating our approach on complex and real-world datasets would be important, publicly available labeled datasets were mostly collected in scripted settings and/or unrealistic environments and lack real-world diversity. Hence, for the sake of this work, we selected two publicly available smart-home datasets that include several types of environmental sensors, with one of them including also inertial sensors from a smartwatch. These two datasets represent different smart home environments and subjects, providing a preliminary indication that our approach generalizes in different conditions. To the best of our knowledge, the most realistic public and labeled datasets are mostly PIR-based deployments, like CASAS (Cook et al., 2009). However, due to noise and limited semantics (as we will discuss in detail in Section 5.3), these datasets pose challenges for LLM-based methods and we excluded them from our analysis. Although public and more realistic unlabeled datasets are available, labeled data is essential for our analysis, as it allows us to demonstrate that wrong predictions are often associated with biased explanations leading to over-reliance (as we discuss in Section 5.1.) Hence, the two public datasets that we consider in our experiments are UCI ADL (Ordónez et al., 2013) and MARBLE (Arrotta et al., 2021).
@@ 4.1.2. MARBLE
The MARBLE (Arrotta et al., 2021) dataset includes data from $12$ subjects in both single and multi-inhabitant scenarios. For the sake of this work, we only consider the single-inhabitant scenarios of the dataset. The data was collected using smart plugs (to monitor stove and TV usage), magnetic (to monitor the interaction with drawers and cabinets), and pressure sensors (to monitor the usage of dining room chairs and the couch in the living room). Additionally, each subject wears a smartwatch to capture inertial data generated by hand gestures. Consistently with the work in (Arrotta et al., 2022a), we only differentiate between static and dynamic gestures. To detect incoming and outgoing phone calls, an Android application was installed on the smartphone of each subject. The original dataset includes 13 activities, but, following (Civitarese et al., 2025), we consider the following $11$ activities: clearing table, eating, entering home, leaving home, phone call, preparing cold meal, preparing hot meal, setting up table, taking medicines, using pc, and watching tv.
@@ 4.2. Baselines
  • ADL-LLM (Civitarese et al., 2025): a recent zero-shot ADL recognition method based on LLMs, shown to be superior to other state-of-the-art smart-home LLM-based approaches. We consider this method as a baseline for the recognition rate, as our LLMe2e approach is an extension and adaptation of it. Differently from LLMe2e, each window is translated into a sentence in natural language with a heuristic approach, rather than providing just the information of sensors’ triggering events like in the JSON template used in LLMe2e. ADL-LLM does not explicitly generate explanations, but we consider it to compare the recognition rate.
@@ 4.3. Experimental Setup
We developed the prototypes of LLMe2e and LLMExplainer in Python, querying the OpenAI ‘gpt-4o’ LLM model through the official OpenAI library, setting the temperature parameter to 0.